An investment project has annual cash inflows of $6,100, $7,200, $8,000 for the next four years, respectively, and $9,300, and a discount rate of 17 percent.

  

What is the discounted payback period for these cash flows if the initial cost is $9,500?

Multiple Choice

  • 3.62 years

  • 0.81 years

  • 1.81 years

  • 1.31 years

  • 2.56 years
Answer #1

Present value of year 1 cash flow = 6100 /(1 + 0.17) = 5,213.6752

Present value of year 2 cash flow = 7200 /(1 + 0.17)2 = 5,259.6976

Present value of year 3 cash flow = 8000 /(1 + 0.17)3 = 4,994.96445

Present value of year 4 cash flow = 9300 /(1 + 0.17)4 = 4,962.94545

Cumulative cash flow for year 0 = -9500

Cumulative cash flow for year 1 = -9500 + 5,213.6752 = -4,286.3248

Cumulative cash flow for year 2 = -4,286.3248 + 5,259.6976 = 973.37

4,286.3248 / 5,259.6976 = 0.81

Discounted payback period = 1 +0.81 = 1.81 years